2. Let P(x):

x  x . (10 pts)

2

Find

a. P (1)

b. P (4)

c.  n, P(n)

d.  n,P(n)


1
Expert's answer
2021-09-09T16:52:07-0400

"a)\\
P(x>-1)= 1-P(\frac{x+2}{\sqrt4}≤\frac{-1+2}{\sqrt4})\\
= 1-P(Z≤\frac{1}{2})= 1- \phi(0.5)\\
=1-0.69146=0.30854\\
b)\\
P(x=0)= P(\frac{x+2}{\sqrt4}≤\frac{0+2}{\sqrt4})\\
=P(Z=1)= 0\\
c)\\
P(x<0)= P(\frac{x+2}{\sqrt4}<\frac{0+2}{\sqrt4})\\
= 1-P(Z<1)=\phi(1)\\
=0.84134\\
d)\\
P(-2<x<0)= P(\frac{-2+2}{\sqrt4}<\frac{x+2}{\sqrt4}<\frac{0+2}{\sqrt4})\\
= P(0<Z<\frac{2}{2})= \phi(1)-\phi(0)\\
=0.84134-0.5\\
=0.3413"